OpenStudy (anonymous):

Solve for x: 2/3 (x - 4) = 2x.

OpenStudy (deadly_roses):

x should equal -2

OpenStudy (anonymous):

Could you explain how you got that?

OpenStudy (tanya123):

(2/3)*(x-4) = 2*x // - 2*x (2/3)*(x-4)-(2*x) = 0 (2/3)*(x-4)-2*x = 0 2/3*(x-4)-2*x = 0 2/3*(x-4)-2*x = 0 -4/3*x-8/3 = 0 -4/3*x-8/3 = 0 -4/3*x-8/3 = 0 // + 8/3 -4/3*x = 8/3 // : -4/3 x = 8/3/(-4/3) x = -2 x = -2

OpenStudy (anonymous):

Thank you!!!

OpenStudy (tanya123):

:*